# Unsupervised Interpretable Basis Extraction for Concept – Based Visual Explanations

### v2

@@ Abstract
An important line of research attempts to explain CNN image classifier predictions and intermediate layer representations in terms of human understandable concepts. In this work, we expand on previous works in the literature that use annotated concept datasets to extract interpretable feature space directions and propose an unsupervised post-hoc method to extract a disentangling interpretable basis by looking for the rotation of the feature space that explains sparse one-hot thresholded transformed representations of pixel activations. We do experimentation with existing popular CNNs and demonstrate the effectiveness of our method in extracting an interpretable basis across network architectures and training datasets. We make extensions to the existing basis interpretability metrics found in the literature and show that, intermediate layer representations become more interpretable when transformed to the bases extracted with our method. Finally, using the basis interpretability metrics, we compare the bases extracted with our method with the bases derived with a supervised approach and find that, in one aspect, the proposed unsupervised approach has a strength that constitutes a limitation of the supervised one and give potential directions for future research.
@@
CNN image classifiers have demonstrated outstanding performance in real-world tasks. They can be used in robotics, visual understanding, automatic risk assessment and more. However, to a human expert, CNNs are often black-boxes and the reasoning behind their predictions can be unclear. Recent advances in explainable and interpretable artificial intelligence (XAI and IAI), attempt to shed light on this process. In an attempt to understand intermediate layer representations, one can project them onto a feature space basis that quantifies the presence of different concepts in the representation. This basis is called interpretable because it can make representations more understandable. In the typical approach, constructing an interpretable basis requires access to annotations. This work proposes a novel unsupervised method to learn such a basis, without the need for explicit labels. This can ease the process of obtaining explanations, eliminate annotation costs, save time, and eventually help humans debug and trust deep models.
@@ 1 Introduction
Figure 2: Left: In a standard convolution layer with $D$ filters, all the filters work together to transform each input patch to a feature vector of spatial dimensionality $1\times 1$. Each spatial element $\mathbf{p}$ of the transformed representation, is assigned a feature vector $\mathbf{x_{p}}\in\mathbb{R}^{D}$ which lies in the co-domain of the transformation function. Thus the dimensionality of the feature space equals the number of filters in the layer, and each spatial element of the transformed representation, constitutes a sample in this feature space. Middle: To find an interpretable basis in the aforementioned feature space in a supervised way, it means to train a set of linear classifiers (concept detectors), one for each interpretable concept, by using feature vectors corresponding to image patches containing the concept. Right: In case classifier training succeeds, the application of the classifier rule ($\mathbf{w}^{T}\mathbf{x_{p}}-b>0$) at each spatial element of the representation $\mathbf{x_{p}}$, produces a binary mask which is active for pixels corresponding to image patches containing the classifier’s concept. We observe, that in a successfully learned interpretable basis, a single pixel is classified positively by at most one classifier, among a group of classifiers that are trained to detect mutually-exclusive concepts.
@@
Despite the impressive performance of convolutional neural networks (CNNs) in computer vision image classification tasks [1, 2, 3, 4], the understanding of their inner workings still remains a challenge. In an attempt to shed light into the CNN “black-box”, the scientific community tries to understand the properties of the intermediate layers’ feature space. Early research [5] showed that any possible direction in this feature space may have a semantic meaning, i.e: feature vectors that maximally activate a direction, correspond to image patches that share some sort of semantic concept. For instance, image patches of car doors, cat heads or people’s faces, maximally activate different directions of this feature space. Beyond this early result, more recently, rigorous experimentation showed that linear separability of features corresponding to different semantic concepts, increases towards the top layer [6]. The latter has been attributed to the top layer’s linearity and the fact that intermediate layers are enforced to produce representations that are helpful to solve the task at hand.
@@
The fact that linear separation of concept representations is possible (especially for layers near the top) [5], has motivated attempts in finding feature space directions for specific concepts [7] and constructing an interpretable feature space basis [8]. In an interpretable basis, each basis vector points towards the direction of a concept’s representations. Projecting a representation onto a basis vector, quantifies the presence of the respective concept in the representation. An interpretable basis can help to obtain possible explanations regarding the CNN and its predictions. When considering the basis vectors as concept embeddings, an interpretable basis can be used to explain the relationship between concepts and filters, similar to what was proposed in [9]. Moreover, it can also be used to interpret predictions of individual examples [8], or used to quantify the class sensitivity of the CNN with respect to a concept [7, 10].
@@
In the typical approach for computing an interpretable basis, the set of interpretable concepts need to be defined in the form of an annotated concept dataset. Using this dataset, one may have access to labels for intermediate CNN representations. These can be subsequently used to find the orientation of the hyperplane that separates the representations of a concept with respect to representations of other concepts [7], [8]. The interpretable basis is constructed by directly using the hyperplane normals as basis vectors. As with any other supervised approach, using an annotated concept dataset to construct an interpretable basis, may increase the fidelity of explanations obtained via that basis. However, this comes at the cost of obtaining the annotations, which is even more prominent when annotations need to be dense (per-pixel) [8]. Additionally, annotated concept datasets are domain-specific, and thus, explaining CNN classifiers for different domains can become even more costly.
@@
The motivation of this work in based on an observation of how an interpretable basis transforms representations. We explain by examples, that the one-hot, thresholded projection of a representation onto an interpretable basis, results into a new, transformed sparse representation. Thus, we propose a method that is able to suggest a feature space basis which satisfies this property found in interpretable bases. In contrast to the typical approach, the proposed method learns the basis directly from the structure of feature space representations, without requiring access to semantic annotations. In that sense, our method can be considered to be unsupervised. However, without annotations, the final suggested basis vectors are not assigned an explicit concept name. In a real-world setting, the concept name associated with a basis vector could be identified by inspecting samples of image patches whose projected representations onto the vector are maximum. For evaluation purposes though, a procedure to label the basis is required, by assigning a concept label to each basis vector, as in [11] or [12].
@@
Our work’s contributions can be summarized to the following: (i) We present a post-hoc, unsupervised method that suggests an interpretable basis for the feature space of a CNN’s intermediate layer. Since post-hoc, the proposed method, applies to pre-trained CNN architectures and does not require any form of retraining them. (ii) Inspired by related work, we propose simple extensions for two basis interpretability metrics. (iii) We provide quantitative evaluation of our method on extracting an interpretable basis for the last layer of popular CNNs, demonstrating applicability to standard architectures. We show that our method is able to improve on the interpretability metrics compared to the interpretability of the natural basis [11], and also compare against a supervised approach [8] to set a baseline for future works and discuss interesting findings that may help future research.
@@ 2 Background & Related Work
In this section we discuss background and related-work in five areas related to this paper. First, we briefly describe prior work on supervised interpretable basis extraction and establish the terminology that is used in this paper. Second, we discuss on supervised and unsupervised discovering of interpretable feature space directions. Third, we highlight how the proposed method differs from other works proposing sparsity as a measure to build inherent-interpretable models. Fourth, we explain the basis labeling problem and potential solutions, and finally, discuss interpretability metrics for assessing the quality of a basis.
@@ 2.1 Supervised interpretable basis extraction and terminology
As already mentioned in the introduction, each basis vector of an interpretable basis points towards the direction of a concept’s representations. To construct an interpretable basis, Zhou et al. [8] trained a set of binary linear classifiers that separate the CNN’s intermediate feature representations based on their semantic meaning. This is accomplished with an densely (per-pixel) annotated concept dataset and implicit use of CNN receptive fields, to assign labels to spatial representation elements of images. Each binary classifier can be considered as a concept detector, since it can separate representations of one concept from representations of other concepts. As already mentioned in Section 1, i) the hyperplane normal directions of the linear classifiers can form a (not necessarily orthogonal, and potentially over/under-complete) basis of the feature space and ii) projecting a representation onto a basis vector, quantifies the presence of the respective concept in the representation. When constructing the basis, each basis vector retains the concept label of it’s respective concept detector. In a strict sense, the concept detector’s bias, which is related to the position of the hyperplane in the feature space, is not part of the basis. For simplicity though, we will retain the association between biases and basis vectors, in such a manner that biases together with the basis vectors form the original concept detectors. For brevity, in this paper, we will use the terms basis, concept detectors, and classifiers almost interchangeably.
@@ 2.2 Discovering concept directions in the CNN feature space
Discovering interpretable directions [5] in the feature space of a CNN image classifier has been previously studied in the literature. In most cases though [6, 8, 7, 10], those directions are directly computed by solving a logistic regression problem that linearly separates CNN’s representations based on their concept label. Thus, these methods rely on the existence of an annotated image dataset. To alleviate the need for concept annotations, Ghorbani et al, [13] proposed a method to automatically group semantically similar image patches out of an unlabeled image dataset. The image patches of each group could then be treated as samples coming from the same concept. Subsequently, the concept samples may be assigned pseudo-labels and can be used as label-representation pairs to reveal each concept’s direction in the feature space of the CNN. The latter may be accomplished via solving the respective logistic regression problem. In contrast to [13], our approach is fundamentally different. In the proposed work, the thresholded projections of CNN’s representations on the learned directions, are sparse. Thus, our work directly tries to exploit existing structure in the CNN’s feature representations, instead of using pseudo-labels to convert the problem to a supervised one.
@@ 2.4 Labeling a feature space basis
We define basis labeling as the procedure of assigning a concept label name to each one of its vectors. When the basis vectors have been learned in a supervised way, the concept label to attribute to the each vector is actually known before learning the vector’s direction. However, when the basis is learned without annotations (such as the current work) or if the natural feature space basis is considered (as in [11] or [12]), attributing meaning to each basis vector requires to put the vector under test. In the testing procedure, each basis vector is accompanied with a (possibly learned) bias (threshold) to form a linear classifier. Then, for all possible concepts, the suitability of the classifier to separate the representations of one concept (positive samples) with respect to the representations of other concepts (negative samples) is evaluated. Finally, each basis vector is assigned a concept label name based on the evaluation metrics of the aforementioned procedure. It is evident, that labeling a basis requires access to a dataset containing concept annotations, such as [11, 17], or [18]. Bau et al. [11] assigned one concept label to each vector of the natural feature space basis based on the Broden dataset (which was also introduced in the same work). Later, Mu et al. [12] used the same dataset to label the natural basis with logical compositions of concepts (e.g. the concept of “blue AND (NOT water)”). In this work, we use [11] to label the bases extracted with our method, while [12] or other potential future works could also be considered.
@@ 3 Motivation
To describe the motivation of our approach, let’s assume that we have access to an interpretable basis of a CNN. Let’s also assume that the basis was successfully learned, i.e: the CNN representations can be linearly separated based on their semantic label. The latter, implies a) the accuracy of the concept detectors is high and b) the CNN has learned to linearly separate (i.e. disentangle [19]) the aforementioned concepts. In b), disentangled representations can be obtained by projecting representations onto the basis. Inversely, in case the CNN representations could not be linearly separated based on their semantic label, it would mean that the accuracy of concept detectors is low and thus concept disentanglement via a linear transformation is not possible.
For example, let’s consider a basis with five concept detectors, one for each element in the set {car, plane, sky, red, white}. Consider the images of the red car and white plane of Fig 2 - right. If we apply the concept detectors to the intermediate representation of an image patch, we observe, that among a group of classifiers detecting mutually-exclusive concepts, only one concept detector classifies the patch positively (i.e. as a patch containing the respective concept). For instance, a patch belonging to the concept car (such as the one located at second row - second column) is not a plane or sky, while it is also red and not white. In that case, {car, plane, sky} is a group of mutually-exclusive concepts, and {red, white} another one. This simple fact, can be summarized to the following observation: projecting a representation to an interpretable basis and thresholding, results into sparse one-hot representations.
In this work, we take a non-standard approach to extract an interpretable basis for the feature space of a CNN. Let us consider a set of linear classifiers. In this case, the classification rule dictates projection on the classifier’s hyperplane normal and hard-thresholding against the classifier’s bias. Based on our previous observation, in case this set of classifiers forms an interpretable basis, applying all the classifiers’ rules to a CNN’s intermediate representation (with hard thresholding) shall result into a new, transformed representation, which is one-hot and sparse. By optimizing basis vectors and biases for this sparsity objective, the proposed method is able to suggest an interpretable basis without requiring an annotated concept dataset.
@@ 4 Proposed Method
Let $D\in\mathbb{N}^{+}$ denote the dimensionality of a layer’s feature space, and $\mathbf{x_{p}}\in\mathbb{R}^{D}$ an element in this space at the spatial location $\mathbf{p}=(x,y)$ (Fig. 2 - left). In a convolutional layer, $D$ equals the layer’s total number of hidden units or, as otherwise mentioned, output channels. Let’s consider a set of $I\in\mathbb{N}^{+},I\leq D$ linear classifiers to form a (possibly) interpretable basis. The $i$-th classifier is characterized by its hyper-plane’s normal direction $\mathbf{w}_{i}\in\mathbb{R}^{D}$ and bias $b_{i}\in\mathbb{R}$, $i\in\mathcal{I},\,\mathcal{I}=\{0,1,...,I-1\}$. Additionally, each one of those classifiers is responsible to quantify the presence of one concept in $\mathbf{x_{p}}$. Last, for the reasons discussed in the Section 5, we also consider $\mathbf{w}_{i}^{T}\mathbf{w}_{i}=1\,\forall i,\,\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0\,\forall\,i,j:i\neq j$, i.e. $\{\mathbf{w}_{i}\}$ should form an orthonormal basis. We consider $I$ to be a hyper-parameter of the method and, without loss of generality, when $I<D$, the orthogonal basis can be trivially completed to dimensionality $D$ in order to represent a rotation of the feature space. The additional $D-I$ directions can be considered as a non-interpretable residual.
@@
To formalize all the previous discussion, consider the standard binary sigmoid classifier $\sigma(\mathbf{w}_{i}^{T}\mathbf{x_{p}}-b_{i})$ which, since $||\mathbf{w}_{i}||=1$ and for full expressivity, requires an additional parameter $M_{i}\in\mathbb{R}^{+}$, such that, $y_{\mathbf{p},i}=\sigma(\frac{1}{M_{i}}(\mathbf{w}_{i}^{T}\mathbf{x_{p}}-b_{i}))$. $M_{i}$ is controlling the margin between the abscissas corresponding to the extremas of the sigmoid and $y_{\mathbf{p},i}\in(0,1)$ denotes the confidence of classifier $i$ to classify $\mathbf{x_{p}}$ positively. Without loss of generality and for mathematical and implementation convenience, we standardize the feature space with batch normalization [20] and without affine parameters. We do so, just after projecting $\mathbf{x_{p}}$ to $\mathbf{w}_{i}$ and before subtracting the bias $b_{i}$ or dividing by $M_{i}$. As already mentioned, the projection of each $\mathbf{x_{p}}$ to the new basis is accomplished via standard $1\times 1$ convolution with $D$ input and $I$ output channels. While searching for $\mathbf{w}_{i}$, the standardization of the feature space allows treating the magnitude of projections $\mathbf{x_{p}}^{T}\mathbf{w}_{i}$, biases $b_{i}$, and margin coefficients $M_{i}$ in the same scale, respectively, regardless of $i$. Thus, this allows us to make a simplification to the parameter space of our model and consider $b_{i}=b$ and $M_{i}=M$ (i.e. equal biases and margins in the standardized space) for all $i$. Orthogonality of the extracted basis is enforced by using [21]. The learnable parameters of our model are simply $\mathbf{w}_{i},b$ and $M$, while $b_{i}$ and $M_{i}$ can be later recovered by inverting the standardization process. A graphical explanation of the principal idea in the 2D feature space is provided in Fig. 4 and the pipeline of the proposed method is given in Figure 5.
@@
[…]

### v3

@@ Abstract
An important line of research attempts to explain CNN image classifier predictions and intermediate layer representations in terms of human-understandable concepts. Previous work supports that deep representations are linearly separable with respect to their concept label, implying that the feature space has directions where intermediate representations may be projected onto, to become more understandable. These directions are called interpretable, and when considered as a set, they may form an interpretable feature space basis. Compared to previous top-down probing approaches which use concept annotations to identify the interpretable directions one at a time, in this work, we take a bottom-up approach, identifying the directions from the structure of the feature space, collectively, without relying on supervision from concept labels. Instead, we learn the directions by optimizing for a sparsity property that holds for any interpretable basis. We experiment with existing popular CNNs and demonstrate the effectiveness of our method in extracting an interpretable basis across network architectures and training datasets. We make extensions to existing basis interpretability metrics and show that intermediate layer representations become more interpretable when transformed with the extracted bases. Finally, we compare the bases extracted with our method with the bases derived with supervision and find that, in one aspect, unsupervised basis extraction has a strength that constitutes a limitation of learning the basis with supervision, and we provide potential directions for future research.
@@
CNN image classifiers have demonstrated outstanding performance in real-world tasks. They can be used in robotics, visual understanding, automatic risk assessment, and more. However, to a human expert, CNNs are often black-boxes and the reasoning behind their predictions can be unclear. Recent advances in explainable and interpretable artificial intelligence (XAI and IAI) attempt to shed light on this process. In an attempt to understand intermediate layer representations, one can project them onto a feature space basis that quantifies the presence of different concepts in the representation. This basis is called interpretable because it can make representations more understandable. In the typical approach, constructing an interpretable basis requires access to annotations. This work proposes a novel unsupervised method to learn such a basis, without the need for explicit labels. This can ease the process of obtaining explanations, eliminate annotation costs, save time, and eventually help humans debug and trust deep models.
@@ 1 Introduction
Figure 2: Left: In a standard convolution layer with $D$ filters, all the filters work together to transform each input patch to a feature vector of spatial dimensionality $1\times 1$. Each spatial element $\mathbf{p}$ of the transformed representation, is assigned a feature vector $\mathbf{x_{p}}\in\mathbb{R}^{D}$ which lies in the co-domain of the transformation function. Thus the dimensionality of the feature space equals the number of filters in the layer, and each spatial element of the transformed representation, constitutes a sample of this feature space. Middle: To find an interpretable basis for the aforementioned feature space in a supervised way, it means to train a set of linear classifiers (concept detectors), one for each interpretable concept, to separate feature vectors corresponding to image patches with and without the concept. Right: In case classifier training succeeds, the application of the classifier rule ($\mathbf{w}^{T}\mathbf{x_{p}}-b>0$) to each spatial element of the representation $\mathbf{x_{p}}$, produces a binary mask which is active for pixels corresponding to image patches containing the concept. We observe, that in a successfully learned interpretable basis, a single pixel is classified positively by at most one classifier, among a group of classifiers that are trained to detect mutually-exclusive concepts.
@@
Despite the impressive performance of convolutional neural networks (CNNs) in computer vision image classification tasks [1, 2, 3, 4], the understanding of their inner workings still remains a challenge. In an attempt to shed light on the CNN “black-box”, the scientific community tries to understand the properties of the intermediate layers’ feature space. Early research [5] showed that any possible direction in this feature space may have a semantic meaning, i.e., feature vectors that maximally activate a direction correspond to image patches that share some sort of semantic concept. For instance, image patches of car doors, cat heads or people’s faces maximally activate different directions of this feature space. Beyond this early result, more recently, rigorous experimentation showed that linear separability of features corresponding to different semantic concepts increases towards the top layer [6]. The latter has been attributed to the top layer’s linearity and the fact that intermediate layers are enforced to produce representations that are helpful to solve the task at hand.
@@
The fact that linear separation of concept representations is possible (especially for layers near the top) [5], has motivated attempts in finding feature space directions for specific concepts [7] and constructing an interpretable feature space basis [8]. In an interpretable basis, each basis vector points towards the direction of a concept’s representations. Projecting a representation onto a basis vector quantifies the presence of the respective concept in the representation. An interpretable basis can help to obtain possible explanations regarding the CNN and its predictions. When considering the basis vectors as concept embeddings, an interpretable basis can be used to explain the relationship between concepts and filters, similar to what was proposed in [9]. Moreover, it can also be used to interpret predictions of individual examples [8], or used to quantify the class sensitivity of the CNN with respect to a concept [7, 10].
@@
In the typical approach for computing an interpretable basis, the set of interpretable concepts needs to be defined in the form of an annotated concept dataset. Using this dataset, one may have access to labels for intermediate CNN representations. These can be subsequently used to find the orientation of the hyperplane that separates the representations of a concept with respect to representations of other concepts [7], [8]. The interpretable basis is constructed by directly using the hyperplane normals as basis vectors. As with any other supervised approach, using an annotated concept dataset to construct an interpretable basis may increase the fidelity of explanations obtained via that basis. However, this comes at the cost of obtaining the annotations, which is even more prominent when annotations need to be dense (per-pixel) [8]. Additionally, annotated concept datasets are domain-specific, and thus, explaining CNN classifiers for different domains can become even more costly.
@@
The motivation of this work is based on an observation of how an interpretable basis transforms representations. We explain by examples that projecting a representation onto an interpretable basis and hard thresholding results in a new, sparse, binary representation. Thus, we propose a method that is able to suggest a feature space basis which satisfies this property that holds for interpretable bases. In contrast to the typical approach, the proposed method learns the basis directly from the structure of feature space representations, without requiring access to semantic annotations. In that sense, our method can be considered to be unsupervised. However, without annotations, the final suggested basis vectors are not assigned an explicit concept name. In a real-world setting, the concept name associated with a basis vector could be identified by inspecting samples of image patches whose projected representations onto the vector are maximum. For evaluation purposes though, a procedure to label the basis is required, by assigning a concept label to each basis vector, as in [11] or [12].
@@
Our work’s contributions can be summarized to the following: (i) We present a post-hoc, unsupervised method that suggests an interpretable basis for the feature space of a CNN’s intermediate layer. Since post-hoc, the proposed method applies to pre-trained CNN architectures and does not require any form of retraining them. (ii) Inspired by related work, we propose simple extensions for two basis interpretability metrics. (iii) We provide a quantitative evaluation of our method on extracting an interpretable basis for the last layer of popular CNNs, demonstrating applicability to standard architectures. We show that our method is able to improve on the interpretability metrics compared to the interpretability of the natural basis [11], and also compare against a supervised approach [8] to set a baseline for future works and discuss interesting findings that may help future research.
@@ 2 Background & Related Work
In this section we discuss background and related work in five areas related to our work. First, we briefly describe prior work on supervised interpretable basis extraction and establish the terminology that is used in this article. Second, we discuss supervised and unsupervised discovering of interpretable feature space directions. Third, we highlight how the proposed method differs from other works proposing sparsity as a measure to build inherently interpretable models. Fourth, we explain the basis labeling problem and potential solutions, and finally, discuss interpretability metrics for assessing the quality of a basis.
@@ 2.1 Supervised interpretable basis extraction and terminology
As already mentioned in the introduction, each basis vector of an interpretable basis points towards the direction of a concept’s representations. To construct an interpretable basis, Zhou et al. [8] trained a set of binary linear classifiers that separate the CNN’s intermediate feature representations based on their semantic meaning. This is accomplished with an densely (per-pixel) annotated concept dataset and implicit use of CNN receptive fields, to assign labels to spatial representation elements of images. Each binary classifier can be considered as a concept detector, since it can separate representations of one concept from representations of other concepts. As already mentioned in Section 1, i) the hyperplane normal directions of the linear classifiers can form a (not necessarily orthogonal, and potentially over/under-complete) basis of the feature space and ii) projecting a representation onto a basis vector quantifies the presence of the respective concept in the representation. When constructing the basis, each basis vector retains the concept label of its respective concept detector. In a strict sense, the concept detector’s bias, which is related to the position of the hyperplane in the feature space, is not part of the basis. For simplicity though, we will retain the association between biases and basis vectors, in such a manner that biases together with the basis vectors form the original concept detectors. For brevity, in this paper, we will use the terms basis, concept detectors, and classifiers almost interchangeably.
@@ 2.2 Discovering concept directions in the CNN feature space
Discovering interpretable directions [5] in the feature space of a CNN image classifier has been previously studied in the literature. In most cases though [6, 8, 7, 10], those directions are directly computed by solving a logistic regression problem that linearly separates CNN’s representations based on their concept label. Thus, these methods rely on the existence of an annotated image dataset. To alleviate the need for concept annotations, Ghorbani et al. [13] proposed a method to automatically group semantically similar image patches out of an unlabeled image dataset. The image patches of each group could then be treated as samples coming from the same concept. Subsequently, the concept samples may be assigned pseudo-labels and can be used as label-representation pairs to reveal each concept’s direction in the feature space of the CNN. The latter may be accomplished via solving the respective logistic regression problem. Ghorbani’s approach, even though automated, remains top-down, automating the process of concept speculation. In contrast to [13], our approach is fundamentally different. Our work directly tries to exploit existing structure in the CNN’s feature representations, instead of using pseudo-labels to convert the problem to a supervised one.
@@ 2.4 Labeling a feature space basis
We define basis labeling as the procedure of assigning a concept label name to each one of its vectors. When the basis vectors have been learned in a supervised way, the concept label to attribute to each vector is actually known before learning the vector’s direction. However, when the basis is learned without annotations (such as the current work) or if the natural feature space basis is considered (as in [11] or [12]), attributing meaning to each basis vector requires putting the vector under test. In the testing procedure, each basis vector is accompanied by a (possibly learned) bias (threshold) to form a linear classifier. Then, for all possible concepts, the suitability of the classifier to separate the representations of one concept (positive samples) with respect to the representations of other concepts (negative samples) is evaluated. Finally, each basis vector is assigned a concept label name based on the evaluation metrics of the aforementioned procedure. It is evident that labeling a basis requires access to a dataset containing concept annotations, such as [11, 17], or [18]. Bau et al. [11] assigned one concept label to each vector of the natural feature space basis based on the Broden dataset (which was also introduced in the same work). Later, Mu et al. [12] used the same dataset to label the natural basis with logical compositions of concepts (e.g. the concept of “blue AND (NOT water)”). In this work, we use [11] to label the bases extracted with our method, while [12] or other potential future works could also be considered.
@@ 3 Motivation
To describe the motivation of our approach, let’s assume that we have access to an interpretable basis of a CNN. Let’s also assume that the basis was successfully learned, i.e., the CNN representations can be linearly separated based on their semantic label. The latter implies a) the accuracy of the concept detectors is high and b) the CNN has learned to linearly separate (i.e. disentangle [19]) the aforementioned concepts. In b), disentangled representations can be obtained by projecting representations onto the basis. Inversely, in case the CNN representations could not be linearly separated based on their semantic label, it would mean that the accuracy of concept detectors is low and thus concept disentanglement via a linear transformation is not possible.
For example, let’s consider a basis with five concept detectors, one for each element in the set {car, plane, sky, red, white}. Consider the images of the red car and white plane of Fig 2 - right. If we apply the concept detectors to the intermediate representation of an image patch, we observe that among a group of classifiers detecting mutually exclusive concepts, only one concept detector classifies the patch positively (i.e. as a patch containing the respective concept). For instance, a patch belonging to the concept car (such as the one located at second row - second column) is not a plane or sky, while it is also red and not white. In that case, {car, plane, sky} is a group of mutually exclusive concepts, and {red, white} another one. This simple fact can be summarized to the following observation: projecting a representation to an interpretable basis and hard thresholding, results into sparse binary representations.
In this work, we take a non-standard approach to extract an interpretable basis for the feature space of a CNN. Let us consider a set of linear classifiers. In this case, the classification rule dictates projection on the classifier’s hyperplane normal and hard-thresholding against the classifier’s bias. Based on our previous observation, in case this set of classifiers forms an interpretable basis, applying all the classifiers’ rules to a CNN’s intermediate representation (with hard thresholding) shall result in a new, transformed representation, which is binary and sparse. By optimizing basis vectors and biases for this sparsity objective, the proposed method is able to suggest an interpretable basis without requiring an annotated concept dataset.
@@ 4 Proposed Method
Let $D\in\mathbb{N}^{+}$ denote the dimensionality of a layer’s feature space, and $\mathbf{x_{p}}\in\mathbb{R}^{D}$ an element in this space at the spatial location $\mathbf{p}=(x,y)$ (Fig. 2 - left). In a convolutional layer, $D$ equals the layer’s total number of hidden units or, as otherwise mentioned, output channels. Let’s consider a set of $I\in\mathbb{N}^{+},I\leq D$ linear classifiers to form a (possibly) interpretable basis. The $i$-th classifier is characterized by its hyper-plane’s normal direction $\mathbf{w}_{i}\in\mathbb{R}^{D}$ and bias $b_{i}\in\mathbb{R}$, $i\in\mathcal{I},\,\mathcal{I}=\{0,1,...,I-1\}$. Additionally, each one of those classifiers is responsible to quantify the presence of one concept in $\mathbf{x_{p}}$. Last, for the reasons discussed in the Section 5, we also consider $\mathbf{w}_{i}^{T}\mathbf{w}_{i}=1\,\forall i,\,\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0\,\forall\,i,j:i\neq j$, i.e. $\{\mathbf{w}_{i}\}$ should form an orthonormal basis. We consider $I$ to be a hyper-parameter of the method and, without loss of generality, when $I<D$, the orthogonal basis can be trivially completed to dimensionality $D$ in order to represent a rotation of the feature space. The additional DID-I directions can be considered as a non-interpretable residual.
@@
To formalize all the previous discussion, consider the standard binary sigmoid classifier $\sigma(\mathbf{w}_{i}^{T}\mathbf{x_{p}}-b_{i})$ which, since $\mathbf{w}_{i}||=1$ and for full expressivity, requires an additional parameter $M_{i}\in\mathbb{R}^{+}$, such that, $y_{\mathbf{p},i}=\sigma(\frac{1}{M_{i}}(\mathbf{w}_{i}^{T}\mathbf{x_{p}}-b_{i}))$. $M_{i}$ is controlling the margin between the abscissas corresponding to the extremas of the sigmoid and $y_{\mathbf{p},i}\in(0,1)$ denotes the confidence of classifier $i$ to classify $\mathbf{x_{p}}$ positively. Without loss of generality and for mathematical and implementation convenience, we standardize the feature space with batch normalization [20] and without affine parameters. We do so, just after projecting $\mathbf{x_{p}}$ to $\mathbf{w}_{i}$ and before subtracting the bias $b_{i}$ or dividing by $M_{i}$. As already mentioned, the projection of each $\mathbf{x_{p}}$ to the new basis is accomplished via standard $1\times 1$ convolution with $D$ input and $I$ output channels. While searching for $\mathbf{w}_{i}$, the standardization of the feature space allows treating the magnitude of projections $\mathbf{x_{p}}^{T}\mathbf{w}_{i}$, biases $b_{i}$, and margin coefficients $M_{i}$ in the same scale, respectively, regardless of $i$. Thus, this allows us to make a simplification to the parameter space of our model and consider $b_{i}=b$ and $M_{i}=M$ (i.e. equal biases and margins in the standardized space) for all $i$. Orthogonality of the extracted basis is enforced by using [21]. The learnable parameters of our model are simply $\mathbf{w}_{i},b$ and $M$, while $b_{i}$ and $M_{i}$ can be later recovered by inverting the standardization process. A graphical explanation of the principal idea in the 2D feature space is provided in Fig. 4 and the pipeline of the proposed method is given in Fig. 5.
@@
with $R=I-\sum_{\mu}\lfloor\alpha_{\mu}I\rfloor$, and $\lfloor\cdot\rfloor$ denoting the floor operation. The previous procedure ensures that $\sum_{\mu}n_{\mu}=I$ while $n_{\mu}$ remains integer. Let $\tau\in[0,1]$ denote a percentage threshold over the total number of pixels in the dataset. We distribute $\tau$ across the concept detectors using a weighting scheme that utilizes the same weight $\omega_{\mu}\in\mathbb{R}^{+}$ for all detectors in the same partition. The $i$-th concept detector is penalized whenever the percentage of pixels assigned to it falls below the threshold $\nu_{i}$ given below: